# SNet: Innovating Stereo Matching and Semantic Segmentation with a Single-Branch Semantic Stereo Network in Satellite Epipolar Imagery

### v1

@@ Abstract
Stereo matching and semantic segmentation are significant tasks in binocular satellite 3D reconstruction. However, previous studies primarily view these as independent parallel tasks, lacking an integrated multitask learning framework. This work introduces a solution, the Single-branch Semantic Stereo Network (SNet), which innovatively combines semantic segmentation and stereo matching using Self-Fuse and Mutual-Fuse modules. Unlike preceding methods that utilize semantic or disparity information independently, our method identifies and leverages the intrinsic link between these two tasks, leading to a more accurate understanding of semantic information and disparity estimation. Comparative testing on the US3D dataset proves the effectiveness of our SNet. Our model improves the mIoU in semantic segmentation from 61.38 to 67.39, and reduces the D1-Error and average endpoint error (EPE) in disparity estimation from 10.051 to 9.579 and 1.439 to 1.403 respectively, surpassing existing competitive methods. Our codes are available at: https://github.com/CVEO/S3Net.
@@ 1 Introduction
In this study, we introduce the end-to-end Single-branch Semantic Stereo Network (SNet), a novel approach that unifies semantic segmentation and disparity estimation to leverage the inherent correlation between semantic content and disparity. In doing so, it captures their inherent connection, thus improving semantic understanding and disparity accuracy. This closely coupled multi-task learning allows for a better understanding of complex scenes, consequently boosting robustness and generalizability.
@@ 2 Methodology
The overall architecture of SNet is shown in Fig.1. Unlike traditional designs, our network uses a single branch configuration. It starts with the Disparity-Classification Spatial Feature Extraction Module (DCSFEM) that extracts features from the left and right images, generating a 4D cost volume containing semantic and disparity information. The Mutual-Fuse Module (MFM) then processes this volume, integrating disparity and semantic information. Finally, subjecting the cost volume to both trilinear and bilinear upsampling strategies results in two outputs at the original resolution: a disparity map and a pixel-level classification map.
@@ 2.2 Cost Volume
Unlike traditional cost volume stacking methods such as PSMNet [4] and SNet [3], we employ a selective approach towards stacking the multi-scale image features from both the left and right images, after they are processed through DCSFEM. The resultant structure forms a 4D cost volume (represented as ) with dimensions corresponding to the height, width, number of disparities, and number of feature maps , which inherently includes an array of rich disparity and semantic features. The topmost layer of disparity in this 4D cost volume is reserved for semantic information, whereas the successive layers encapsulate disparity information from multi-scale features.
@@ 3.3.1 Compared Methods
Our analysis primarily involves two different aspects: For the purpose of disparity estimation comparison, we conduct an exhaustive evaluation of currently superior algorithms such as PSMNet [4], GwcNet [5], GANet [6], CFNet [7], and SNet [3]; In evaluating the task of semantic segmentation, we have selected advanced segmentation algorithms including SegFormer [8], PSPNet [9], SDFCNv2 [10], and HRNetV2 [11].
@@ 3.3.2 Stereo Matching task
As shown in Table 2, our proposed SNet significantly outperformed other methods, demonstrating lower D1-Error (9.579) and EPE (1.403). As shown in Fig.2, although PSMNet and SNet both showed good results, our method presented more detailed and accurate disparity details, especially at object edges and in areas with rich textures. As shown in the red box in the Fig.2, our method better reflected the outline of the building and the edge information of the water.
| Methods | PSMNet | GwcNet | GANet | CFNet | SNet | Ours |
@@ 3.3.3 Semantic Segmentation task
According to the Table 3, our SNet demonstrates outstanding performance across different categories and performs better in specific scenarios (such as Water and Bridge). As shown in Fig.3, although the PSPNet and HRNetV2 respectively show good results on water bodies and buildings, our method presents clearer contours for categories such as buildings, trees, and water.
@@ 4 Conclusion
In the research, we introduce a novel multitask learning framework called the (SNet) to simultaneously infer disparity maps and classification maps. The uniqueness of our method stems from capitalizing on the strong correlation between these tasks, effectively integrating them via self-fusion and mutual fusion modules for mutual enhancement. Notably, the evaluation results obtained from the US3D dataset and the comparison with other models affirm the feasibility and exceptional performance of our task framework. In the future, we hope to extend the results of this study to applications in multiview stereo matching and 3D reconstruction of multi-sensor data.

### v3

# S3Net: Innovating Stereo Matching and Semantic Segmentation with a Single-Branch Semantic Stereo Network in Satellite Epipolar Imagery

@@ Abstract
Stereo matching and semantic segmentation are significant tasks in binocular satellite 3D reconstruction. However, previous studies primarily view these as independent parallel tasks, lacking an integrated multitask learning framework. This work introduces a solution, the Single-branch Semantic Stereo Network (S3Net), which innovatively combines semantic segmentation and stereo matching using Self-Fuse and Mutual-Fuse modules. Unlike preceding methods that utilize semantic or disparity information independently, our method identifies and leverages the intrinsic link between these two tasks, leading to a more accurate understanding of semantic information and disparity estimation. Comparative testing on the US3D dataset proves the effectiveness of our S3Net. Our model improves the mIoU in semantic segmentation from 61.38 to 67.39, and reduces the D1-Error and average endpoint error (EPE) in disparity estimation from 10.051 to 9.579 and 1.439 to 1.403 respectively, surpassing existing competitive methods. Our codes are available at: https://github.com/CVEO/S3Net.
@@ 1 Introduction
In this study, we introduce the end-to-end Single-branch Semantic Stereo Network (S3Net), a novel approach that unifies semantic segmentation and disparity estimation to leverage the inherent correlation between semantic content and disparity. In doing so, it captures their inherent connection, thus improving semantic understanding and disparity accuracy. This closely coupled multi-task learning allows for a better understanding of complex scenes, consequently boosting robustness and generalizability.
@@ 2 Methodology
The overall architecture of S3Net is shown in Fig.1. Unlike traditional designs, our network uses a single branch configuration. It starts with the Disparity-Classification Spatial Feature Extraction Module (DCSFEM) that extracts features from the left and right images, generating a 4D cost volume containing semantic and disparity information. The Mutual-Fuse Module (MFM) then processes this volume, integrating disparity and semantic information. Finally, subjecting the cost volume to both trilinear and bilinear upsampling strategies results in two outputs at the original resolution: a disparity map and a pixel-level classification map.
@@ 2.2 Cost Volume
Unlike traditional cost volume stacking methods such as PSMNet [4] and S2Net [3], we employ a selective approach towards stacking the multi-scale image features from both the left and right images, after they are processed through DCSFEM. The resultant structure forms a 4D cost volume (represented as ) with dimensions corresponding to the height, width, number of disparities, and number of feature maps , which inherently includes an array of rich disparity and semantic features. The topmost layer of disparity in this 4D cost volume is reserved for semantic information, whereas the successive layers encapsulate disparity information from multi-scale features.
@@ 3.3.1 Compared Methods
Our analysis primarily involves two different aspects: For the purpose of disparity estimation comparison, we conduct an exhaustive evaluation of currently superior algorithms such as PSMNet [4], GwcNet [5], GANet [6], CFNet [7], and S2Net [3]; In evaluating the task of semantic segmentation, we have selected advanced segmentation algorithms including SegFormer [8], PSPNet [9], SDFCNv2 [10], and HRNetV2 [11].
@@ 3.3.2 Stereo Matching task
As shown in Table 2, our proposed S3Net significantly outperformed other methods, demonstrating lower D1-Error (9.579) and EPE (1.403). As shown in Fig.2, although PSMNet and S2Net both showed good results, our method presented more detailed and accurate disparity details, especially at object edges and in areas with rich textures. As shown in the red box in the Fig.2, our method better reflected the outline of the building and the edge information of the water.
| Methods | PSMNet | GwcNet | GANet | CFNet | S2Net | Ours |
@@ 3.3.3 Semantic Segmentation task
According to the Table 3, our S3Net demonstrates outstanding performance across different categories and performs better in specific scenarios (such as Water and Bridge). As shown in Fig.3, although the PSPNet and HRNetV2 respectively show good results on water bodies and buildings, our method presents clearer contours for categories such as buildings, trees, and water.
@@ 4 Conclusion
In the research, we introduce a novel multitask learning framework called the (S3Net) to simultaneously infer disparity maps and classification maps. The uniqueness of our method stems from capitalizing on the strong correlation between these tasks, effectively integrating them via self-fusion and mutual fusion modules for mutual enhancement. Notably, the evaluation results obtained from the US3D dataset and the comparison with other models affirm the feasibility and exceptional performance of our task framework. In the future, we hope to extend the results of this study to applications in multiview stereo matching and 3D reconstruction of multi-sensor data, and further expand the experimentation of this method in various imagery scenarios.